CR12:

The graph below shows the price of different numbers of cans of beans at a store: Which equation can be used to determine p, the cost of b cans of beans? p = 9.50 + b p = 9.50b b = 9.50 + p b = 9.50p

CR12:

1 attachment
CR12:

hi

AZ:

Hey there!

CR12:

can u help me?

AZ:

Do you have an inkling on how to go about this question?

CR12:

uhh i have no idea what inkling means

AZ:

haha, I meant do you have any idea how to begin answering it? Were you able to eliminate any answer choices yet?

CR12:

oh no not yet

AZ:

That's alright. Do you know the slope intercept form of an equation?

CR12:

do you mean if i know if its proportional?

AZ:

no, have you seen y = mx + b where m = slope and b is the y-intercept

CR12:

kinda but more in a lower grade way if u know what i mean

CR12:

i know its like Y = Y x X

AZ:

How are the x and y axises labelled in your graph? Would b, the cost of beans, be x or y?

CR12:

i mean Y = K x X

CR12:

p = Y b = x

AZ:

y = kx is for proportionality, yes. And this particular equation does follow that pattern but not all equations will follow that pattern.

CR12:

ok

CR12:

did u see the screen shot of the graph

AZ:

That's good! We've made some progress now. We need to use the y = mx + b form of an equation y = mx + c You just said that p is our y and b is our x so let's replace y with p and x with b We get p = mb +c

AZ:

Yes, I did. Do you know what the y-intercept is?

CR12:

no im in 7th lol

CR12:

prob should of told u that earlier lol

AZ:

No, that's alright! So the y-intercept is where the line crosses the y-axis. It's basically when x = 0. What is the y-value when x = 0? Just look at the graph

CR12:

ok

CR12:

anyway doesnt that we can eliminate b and d

AZ:

no no not yet! What answer choices can we eliminate first? Remember we said that p was our y so the equation should start with p =

CR12:

76?? i think

AZ:

woah where did you get the number 76 from

CR12:

im lost should we start over

CR12:

because all these big words that i dont know what they mean r confusing me

AZ:

aww, I'm sorry

AZ:

So we agree that p is our y and b is our x right?

CR12:

no im sorry im just waisting ur time

CR12:

yea

AZ:

now you know that equations are generally written as y = mx + c or how you even said y = kx so if p is our y. Then that means we switch the y with a p p = mx + c and we said b was our x. Just switch the x with a b p = mb + c now we have to find out what `m` and `c` are but first we can eliminate a few answer choices which answer choices don't start off with a p?

CR12:

c and d

AZ:

good!

AZ:

Now, I know you haven't learned about y-intercepts. But have you learned about calculating the slope of a line? Does `rise over run` sound familiar to you?

CR12:

no 🙁

AZ:

I wonder how they expect you to answer this mhmm

AZ:

So y-intercept is when our line crosses the y-axis You know what the y-axis is, right?

CR12:

yes

AZ:



CR12:

ok

AZ:

The y-intercept is when our line crosses the y-axis. And it crosses the y-axis when x = 0. So look back at your graph. What is the y-value when the line crosses the y-axis

CR12:

its 0?

AZ:

Yes!

CR12:

YAY

AZ:

Now remember what I said earlier Equations are in the form of y = mx + c that letter c is our y-intercept and if you said it's 0, then that means all we're left with is y = mx (because y = mx + 0 is the same thing as y = mx)

AZ:

And obviously we should replace the x and y with our p and b so we should have something like p = mb

CR12:

ok im following

AZ:

We don't need to calculate the slope (unless you want to) because we can eliminate one more answer without having to solve for what number that 'm' (or slope) is

CR12:

k

AZ:

So between p = 9.50 + b p = 9.50b Knowing that our equation should be something like p = mb where m is some number for the slope what do you eliminate and what's the final answer?

CR12:

elim A?

AZ:

yay! you got it!!

CR12:

yessssss

CR12:

ty ty TYYYYYYYYYY

AZ:

It was my pleasure!! Thank you for being a good sport!

CR12:

thanks for sticking around 🤩

CR12:

btw i faned u 🙂

AZ:

Anytime!